# Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training

### v6

@@ 4.2 Numerical Results
Differentiability Not Enforced $1.64\times 10^{-7}$ $3.20\times 10^{-6}$ $1.02\times 10^{-5}$ $3.73\times 10^{-5}$
Differentiability Enforced $\mathbf{7.86\times 10^{-8}}$ $\mathbf{8.86\times 10^{-7}}$ $\mathbf{5.27\times 10^{-7}}$ $\mathbf{7.87\times 10^{-6}}$
@@
Differentiability Not Enforced $\mathbf{4.41\times 10^{-8}}$ $1.49\times 10^{-7}$ $1.47\times 10^{-5}$ $3.81\times 10^{-4}$
Differentiability Enforced $5.06\times 10^{-8}$ $\mathbf{6.82\times 10^{-8}}$ $\mathbf{2.21\times 10^{-7}}$ $\mathbf{8.42\times 10^{-7}}$
@@
Differentiable Pretraining $\mathbf{5.65\times 10^{-6}}$ $\mathbf{6.53\times 10^{-4}}$ $\mathbf{7.92\times 10^{-7}}$ $\mathbf{5.09\times 10^{-6}}$
Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends. First, the worst performing networks are the Default Networks that rely on randomized (Kaiming) initialization (See also Figure 5). Even the networks that forgo pretraining benefit from initializing with many activation regions. When pretraining constraints are used, they are able to steer gradient descent to the best solutions, resulting in reductions in minimum error of three orders of magnitude over default networks. Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during subsequent unassisted gradient descent.
The last trend to observe is the poor average performance of default networks. In a typical run of these experiments, around half of the default networks collapse. RAAI is able to eliminate most, but not all of the dying ReLU instances due to its probabilistic nature, so it, too, has high mean error.
@@
Our second set of results is shown in Table 3 and in Figure 4. Here the most important impact of utilizing exponentially many linear regions is demonstrated. Not only can more accurate representations of training data be learned, but more linear regions allow the network to better capture underlying nonlinearity to enhance its predictive power in regression tasks. This result is especially significant because it indicates that even in cases where there are fewer data points than linear regions, having the additional regions can still provide performance advantages.
Figure 5: Neuron outputs of a default (Kaiming-initialized) network (left) versus a pretrained variant of our network (right). The reason our network succeeds in using so many linear regions, when a random one fails, is that the number of regions is not a “local” property of parameter space. The first 3 layers of the default network are not adding any new linear regions, and infinitesimal adjustments to the slopes or biases of the lines shown would not change that. In our network, the pretraining guides it to a region of parameter space where $2^{d}$ activation regions are generated. Since the region count is not a local property, we still have a neighborhood of parameter space we can optimize in and maintain efficient neuron use.
@@ 5 Extension to Non-Convex Functions and Higher Dimensions
Figure 6: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $4\times 10^{-7}$. A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$. Notice though that the individual neuron outputs are much messier than in our convex case (see Figure 5) This is because naive gradient descent has all the weights in between the two pretrained component networks to learn unassisted.
The results presented so far are limited to one-dimensional convex functions, but our method can easily be extended. Our constructed networks can be applied to approximate higher-dimensional functions by using these networks as activation functions in a larger network (see Figure 8 in the Appendix for a schematic). For instance, in two dimensions, we take two or more copies of our network and estimate a target function as a linear combination of the networks’ outputs. Because the weights of the linear combination are randomized, each copy of the network will tend to learn different lower-dimensional aspects of the target function, whereby the combination can accurately capture the higher-dimensional function.
The same approach can address non-convex functions since many non-convex functions (including every continuous function with a bounded second derivative (Hessian)) can be expressed as the difference of two convex functions (see Zhang et al. (2018), which also describes how every ReLU network can be decomposed into a difference of two piecewise linear convex functions). We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising. In fact, we highly encourage reviewing Appendix A.10, which demonstrates our method on real-world regression datasets with up to eight input dimensions, as well as image classification problems.
We note that the approach in this section bears similarity to the recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024), which use spline functions as activations of a network where every weight is 1. This is based on the Kolomogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957, 1959), which gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure with one-dimensional activation functions. But by restricting the activation functions to splines, the KAN abandons its theoretical guarantees. Using our networks as activations, however, maintains the theoretical ability to perform polynomial interpolation since they retain the ability to approximate $x^{2}$, and the multiplication gate discussed in Section 2.1 is based on a linear combination of squared terms.
Figures 6 and 7 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks. Rather than directly applying our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect). Two slight modifications are made to each of the component networks to better interact with ReLU. First, two neurons are used for the sum so that a negative component may be maintained in the case of $x^{3}-x$, which would otherwise be zeroed out by ReLU. Second, the triangular functions are generated upside-down, preventing ReLU from zeroing the function outside of the approximation region.
For both $x^{3}-x$ and $r^{3}$, we substantially outperform standard initialization and training procedures. Especially striking is the increased number of linear regions in Figure 7. Visible in Figure 6 is the gradient descent-induced destructuring discussed in Section 4.2. This likely occurs due to the block diagonal structure of the larger weight matrices, which results in many zeros for the second pass of gradient descent to fill in unassisted. This could be addressed either by limiting training only to the block diagonal entries, or by deriving a more advanced pretraining method that makes use of the full matrix. Note that since $x^{3}-x$ is twice differentiable we can write it as $(x^{3}-x+3x^{2})-3x^{2}$, a difference of functions convex on $[-1,1]$ obtained by finding a parabola based on trying to make the derivative of $x^{3}-x$ monotone. Note that this is not a unique decomposition.
@@ 6 Concluding Remarks
This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of efficiently learning nonlinear functions; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased. Our results showed improvements of one to several orders of magnitude in using our initialization and pretraining strategy to train ReLU networks to learn various nonlinear functions, including non-convex and multi-dimensional functions. This finding is particularly powerful since random initialization and gradient descent are not likely to produce an efficient solution on their own, even if it can be proven to exist in the set of sufficiently sized ReLU networks. Although we proposed one preliminary strategy for extending our network for higher-dimensional functions, this strategy has limitations, such as leaving many weights to be filled in by gradient descent (which it generally does not do effectively). We anticipate continuing investigating strategies for improving the generalizability of our work, which we believe is critically important since it offers the possibility of an exponentially more accurate drop-in replacement of linear layers in any architecture. We note that there are other metrics for evaluating a neural network’s expressivity (e.g., Raghu et al. (2017)), which would be interesting to measure for our networks in future work. Overall, we are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU and other classes of neural networks.
Appendix A Appendix
@@ A.1 Initialization and Pretraining Algorithm
The initialization step of our algorithm is to generate a vector $A=[a_{0},a_{1},...a_{n}]^{T}$, where each $a_{i}$ is randomized in $(0,1)$. Given this $A$, the pretraining step of our algorithm sets the weights of the input ($I$), hidden ($H_{i}$, $11\leq i\leq n-$ 1), and output ($O$) layers of the network as follows:
@@
$H_{i}(x)=\begin{bmatrix}1&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&-S_{i}a_{i+1}\\ 0&0&0&S_{i}\end{bmatrix}\times ReLU(H_{i-1}(x))$
$O(x)=\begin{bmatrix}1&S_{n}/a_{n}&-S_{n}/(a_{n}-a_{n}^{2})&0\end{bmatrix}\times ReLU(H_{n-1}(x))$
where $S_{i}$ can either be chosen independently, or chosen based on $A$. In the latter case, Equation 4 gives $S_{i}=s_{i}/s_{i-1}=(1-a_{i})a_{i+1}$. This assignment of $I$, $H$, and $O$ is used in each iteration of the pretraining algorithm.
@@
The network weights can then be initialized once more based on the learned vector $A$, to then update the weights via regular gradient descent. In our experiments, both phases of training (pretraining and gradient descent) ran for 1000 epochs. Full pseudocode is listed in Algorithm 1.
@@ A.2 Terminology
Before presenting further formal mathematical details of our method, we first briefly review a few pieces of basic ReLU network terminology used in the paper. The reader is referred to the Chmielewski-Anders (2020) for an excellent alternative presentation of these terms.
@@ A.3 Mathematical Results
$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x)$
The goal of this section will be to show how to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined and continuous on all of $[0,1]$. We begin by assuming that
$F^{\prime}(x)=\sum_{i=0}^{\infty}s_{i}W^{\prime}_{i}(x)$
We will see that the resulting choice of $s_{i}$ ensures uniform convergence of the derivative terms, so that the derivative of the infinite sum is indeed infinite sum of the derivatives. Fortunately, the left and right derivatives $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ already exist everywhere, since each bend in each $W_{i}$ (where the full derivative of $F$ is undefined) is assigned the slope of the line segment to its left or right respectively. The $s_{i}$ scaling values will have to be chosen appropriately so that $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ are equal for all bend points.
@@
Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$. We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends. $B_{i}:=P_{i}\cup V_{i}$. $f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$. The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network. This odd split around layer $i$ makes the proofs cleaner. Figure 9 highlights some important properties about composing triangle functions. Peaks alternate with valleys. Peak locations in one layer become valleys in the next. Valleys in one layer remain valleys in all future layers since 0 is a fixed point of each $T_{i}$. To produce $W_{i}$, each line segment of $W_{i-1}$ becomes a dilated copy of $T_{i}$. Each triangle function has two distinct slopes $1/a_{i}$ and $-1/(1-a_{i})$ which are dilated by the chain rule during the composition. On negative slopes of $W_{i-1}$, the input to layer $i$ is reversed, so those copies of $T_{i}$ are reflected. Due to the reflection, the slopes of $W_{i}$ on each side of a peak or valley are proportional. Alternatively, one could consider that on each side of a peak in $W_{i-1}$, there is a neighborhood of points that are greater than $a_{i}$, and are composed with the same line segment of $T_{i}$ that has slope $-1/(1-a_{i})$. Either way, it’s important to note that the slopes on each side of a bend scale identically during each composition.
@@
Let $x_{k}$ be some point in $P_{i}$, and let $k$ be its index in any list it appears in. To calculate the value of $E_{+}^{\prime}(x_{k})=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x_{k})$, we will have to find the slope of the linear intervals to the immediate right of $x_{k}$ for all $W_{i}$. We will use $R_{x}$ to represent $W_{i+}^{\prime}(x_{k})$. The first term in the sum will be $R_{x}s_{i}$. Since the derivatives of composed functions will multiply from the chain rule, so the value of the next term is $W^{\prime}_{i+1}(x_{k})=T^{\prime}_{i+1}(W_{i}(x_{k}))R_{x}$. $T_{i+1}$ has two linear segments, giving two slope possibilities to multiply by. The correct one to choose is $-1/(1-a_{i+1})$ because it’s ‘active’ around $x_{k}$ ($x_{k}$ is a peak of $W_{i}$, so $W_{i}(x_{k})>a_{i+1}$ for $x\in(B_{i+1}[k-1],B_{i+1}[k+1])$). This gives $W^{\prime}_{i+1}(x_{k})=-R_{x}\frac{s_{i+1}}{(1-a_{i+1})}$. Note that the second term has the opposite sign as the first.
For all remaining terms, since $x_{k}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$. For $x\in(B_{j+1}[k-1],B_{j+1}[k+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the slope $1/a_{j+1}$. Since this slope is positive, every remaining term continues to have the opposite sign as the first term. Summing up all the terms with the coefficients $s_{i}$, and factoring out $R_{x}$ will yield the desired formula. Note that this same derivation applies to the left sided derivatives as well because the ‘active’ slopes of $T_{i+1}(W_{i})$ are all the same whether a bend in $W_{i}$ is approached from the left or the right. The initial slope constant $L_{x}$ will just be different. ∎
@@
$F^{\prime}(x)=f^{\prime}_{i}(x)=\sum_{j=0}^{i-1}s_{j}W^{\prime}_{j}(x)$ (6)
@@
We now prove our main theorem, which shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting approximation converges to a continuously differentiable function. The idea of the proof is that much of the formula for $E^{\prime}(x)$ will be shared between two successive generations of peaks. Once they are both valleys, they will be treated the same by the remaining compositions, so the sizes of their remaining discontinuities will need to be proportional.
Theorem (3.1).
$F^{\prime}(x)$ is continuous on $[0,1]$ only if the scaling coefficients are selected based on $a_{i}$ according to:
$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}$
@@
Figure 10: The derivatives of the first few stages of approximation. Notice that each time a constant segment “splits”, the two neighboring segments adjacent to the split monotonically converge back to the original value (marked with a colored point corresponding to the step it’s from)

### v7

@@ 4.2 Numerical Results
Pretraining (Unregularized) $1.64\times 10^{-7}$ $3.20\times 10^{-6}$ $1.02\times 10^{-5}$ $3.73\times 10^{-5}$
Pretraining (with Thm. 3.1) $\mathbf{7.86\times 10^{-8}}$ $\mathbf{8.86\times 10^{-7}}$ $\mathbf{5.27\times 10^{-7}}$ $\mathbf{7.87\times 10^{-6}}$
@@
Pretraining (Unregularized) $\mathbf{4.41\times 10^{-8}}$ $1.49\times 10^{-7}$ $1.47\times 10^{-5}$ $3.81\times 10^{-4}$
Pretraining (with Thm. 3.1) $5.06\times 10^{-8}$ $\mathbf{6.82\times 10^{-8}}$ $\mathbf{2.21\times 10^{-7}}$ $\mathbf{8.42\times 10^{-7}}$
@@
Pretraining (with Thm. 3.1) $\mathbf{5.65\times 10^{-6}}$ $\mathbf{6.53\times 10^{-4}}$ $\mathbf{7.92\times 10^{-7}}$ $\mathbf{5.09\times 10^{-6}}$
Our first set of results are shown in Tables 1 and 2, wherein we observe several important trends. First, the worst performing networks are the “default networks” that rely on randomized (Kaiming) initialization (see also Figure 5). Even the networks that forgo pretraining benefit from initializing with many activation regions. When pretraining constraints are used, they are able to steer gradient descent to the best solutions, resulting in reductions in minimum error of three orders of magnitude over default networks. Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during subsequent unassisted gradient descent.
The last trend to observe is the poor average performance of default networks. In a typical run of these experiments, around half of the default networks collapse from the dying ReLU phenomenon. RAAI is able to eliminate most, but not all of the dying ReLU instances due to its probabilistic nature, so it, too, has high mean error. This is why we have focused on presenting the minimum error over all the trials.
@@
Our second set of results is shown in Table 3 and in Figure 4. Here the most important impact of utilizing exponentially many linear regions is demonstrated. Not only can more accurate representations of training data be learned, but maintaining more linear regions allows the network to better capture underlying nonlinearity to enhance its predictive power in regression tasks. This result is especially significant because it indicates that even in cases where there are fewer data points than linear regions, having the additional regions can still provide performance advantages.
Figure 5: Neuron outputs of a default (Kaiming-initialized) network (left) versus a pretrained variant of our network (right). Notice that the first two layers of the default network introduce no linear regions - none of the lines cross zero. Any infinitesimal adjustment to the slopes or biases of the lines would not make such an intersection occur. Therefore, the number of linear regions generated by the network cannot be a local property, and we can expect gradient-based optimization to struggle at maximizing the linear region count. Our method uses this non-locality to our advantage. The pretraining phase finds a low-loss solution where $2^{d}$ linear regions are generated, which guarantees a neighborhood of parameter space where $2^{d}$ regions can be maintained while training in the standard parameterization.
@@ 5 Extension to Non-Convex Functions and Higher Dimensions
Figure 6: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $5.52\times 10^{-7}$. A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$.
The results presented so far are limited to one-dimensional convex functions, but our method can easily be extended. Since our constructed networks are one-dimensional, they can be used as the activation functions of a larger network (see Figure 9 in the Appendix for a schematic). Our experimental results show that even using just two copies of our network in this manner can go a long way towards increasing the expressive ability of the technique. For example, two of our networks could be used to build a two-dimensional convex function. This can be done by using one network to build a convex function oriented in the $x$ direction, and the other in the $y$ direction. A positive linear combination of their outputs then yields a two-dimensional convex function. Similarly, a non-convex function can be realized by taking a difference of the outputs of two of our networks. Of course, this pattern can be generalized to an arbitrary-width layer using $n$ of our networks.
Figures 6 and 7 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional convex function $z=r^{3}$ on $[-1,1]$, each by using two of our constructed networks with 5 hidden layers. For both $x^{3}-x$ and $r^{3}$, we substantially outperform standard initialization and training procedures. Especially striking is the increased number of linear regions in Figure 7. Like the previous experiments, these networks were trained using Adam for 1000 epochs at a learning rate of 0.001, and the minimum over 30 trainings was taken. The parameterization switch was made halfway through training. These experiments use a slightly improved pretraining construction (see Appendix A.1): the input domain is modified to be $[-1,1]$, and the network is adjusted so that the $t_{1}$ and $t_{2}$ neurons compose “v” shapes rather than triangles, which allows the network to extrapolate better outside $[-1,1]$. With multiple copies of the network, we implement the calculation of the $d$-th hidden layer as one large combined matrix. Since the hidden neurons of each subnetwork do not connect, the combined matrix takes on a sparse block diagonal structure with a block size of 4. When we switch parameterizations, we release the block diagonal constraint and allow gradient descent to fill in the zeroed weights.
We make a few notes on our preliminary extended technique. First, the class of non-convex functions that can be represented as a difference of convex functions is broad. Zhang et al. (2018) gives an iterative process to decompose an arbitrary ReLU network into a difference of piecewise linear convex functions. Additionally, all functions with bounded second derivatives (or bounded eigenvalues of their Hessian matrix) can be expressed as a difference of two convex functions. Finally, we note that these types of decompositions are not necessarily unique; for example, while $x^{3}-x$ can be expressed as $(x^{3}-x+3x^{2})-3x^{2}$, $3x^{2}$ is merely the minimal parabola needed to make the second derivative positive on $(-1,1)$, and larger coefficients could also be used.
Even though our original technique is not naturally designed for multivariate functions and has to be extended by incorporating it as an activation function in a larger network, we note that doing so has some theoretical backing. The multiplication gate and polynomial interpolation network discussed in the related work (Section 2.1) rely on linear combinations of squared terms. Since our networks can produce $x^{2}$, an arbitrary width and depth arrangement of our networks would retain the ability to perform polynomial interpolation. Secondly, the Kolmogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957, 1959) gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure that is built using only addition and arbitrary continuous one-dimensional activation functions. Essentially, it states that a technique that only works for univariate functions can still be sufficient for representing multivariate functions when deployed in a larger network. The recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024) take their inspiration from this theorem, extending their networks to arbitrary width and depth, but restricting the activations to splines. Even though this does not strictly follow the theorem, they still find empirical success. Similarly, in the case of this work, the functions that the triangular parameterization can directly represent are not dense, as they require a certain degree of self-similarity. Nonetheless, when assembled into a larger network, this may not pose a serious problem.
@@
6 Preliminary Image Classification Results
Since our extended technique can be used to replace any dense layers in arbitrary architectures, we can use it in contexts beyond regression. In our final experiment, we replace the densely connected classifier of VGG-16 (Simonyan & Zisserman, 2014), and then retrain the network on ImageNet (Deng et al., 2009). We initialize with PyTorch’s pretrained convolutional weights both to speed up training and to better isolate the effects of modification to the dense layers. Since VGG-16’s dense layers have a width of 4096, we decided to leave the matrices corresponding to our subnetworks in block diagonal format, rather than allowing them to become dense. This way, our technique only adds a negligible amount of extra operations (roughly $0.5\%$) and does not leave the majority of weights to be filled in by unassisted gradient descent.
Figure 8: Retraining the classifier of a VGG-16 model on ImageNet. The ordinary network uses a learning rate of $1\times 10^{-4}$, and our experimental network uses $1\times 10^{-5}$. At epoch 10, both learning rates decay by a factor of 10, and we switch the parameterization in the experimental network. Both networks retrain slightly higher than PyTorch’s original reported accuracy of $73.3\%$.
Figure 8 shows that our network has an advantage early in training, but the fully trained networks achieve comparable accuracy. This is perhaps unsurprising: ordinarily initialized and trained networks are already extremely good at classification problems despite the pathological properties of random initialization. This result suggests that highly precise representations of decision boundaries may not be critical to performance for such classification problems. In support of this, we noticed that it did not seem to matter if we switched parameterizations with our technique, even though doing so was critical to our performance on the 1D regression problems. A practical-scale regression task may be needed for our method to display a significant advantage; we plan such examples as future work.
@@ 7 Concluding Remarks
This paper focused on exploiting the potential computational complexity advantages neural networks offer for the problem of efficiently learning nonlinear functions; in particular, compelling ReLU networks to approximate functions with exponential accuracy as network depth is linearly increased. Our results showed improvements of one to several orders of magnitude in using our initialization and pretraining strategy to train ReLU networks to learn various nonlinear functions, including non-convex and multi-dimensional functions. This finding is particularly powerful since random initialization and gradient descent are not likely to produce an efficient solution on their own, even if it can be proven to exist in the set of sufficiently sized ReLU networks.
We anticipate the continued development of theoretical improvements to this work. Of particular importance, we think, is either finding a dense set of one-dimensional functions we can represent efficiently with deep networks, or finding a method that more naturally represents multidimensional functions. While works like the KAN show it is possible to perform well without those things, we believe them to be important. The fact that our current technique is limited to creating sparse block diagonal matrices is indicative of a need to determine how to use all the weights of dense layers. Another interesting avenue for increasing the technique’s expressiveness could be to consider the generation of fractals. We disallow such behavior by using Theorem 3.1 to enforce differentiability in the hopes of regularization, but there may be more sophisticated approaches to function modeling that can generate fractals in a controlled manner. Given that the technique in its current form is flexible enough to go into any architecture, it is also important to explore possible use cases. We have shown that it will likely not lead to significant improvements on classification tasks, but the technique may shine when given an appropriate utility-scale regression task. Overall, we are hopeful that future works by our group and others will help illuminate a complete theory for harnessing the potential exponential power of depth in ReLU and other classes of neural networks.
Impact Statement
This paper presents work whose goal is to enable more efficient neural networks. Future advances along this line could enable the use of much smaller networks in practical applications, which could substantially mitigate the rapidly growing issue of energy usage in large learning systems.
Acknowledgments
M.M. would like to thank Xander Neuwirth for insightful conversations. D.H. was supported in part by the U.S. National Science Foundation under award number 2442853. This work was authored in part by the National Renewable Energy Laboratory (NREL), operated by Alliance for Sustainable Energy, LLC, for the U.S. Department of Energy (DOE) under Contract No. DE-AC36-08GO28308. This work was supported by the Laboratory Directed Research and Development (LDRD) Program at NREL. The views expressed in the article do not necessarily represent the views of the DOE or the U.S. Government. The U.S. Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S. Government purposes.
Appendix A Algorithm and Theory
@@ A.1 Initialization and Pretraining Algorithm
The initialization step of our algorithm is to generate a vector $A=[a_{0},a_{1},...a_{n}]^{T}$, where each $a_{i}$ is randomized in $(0,1)$. Given this $A$, the pretraining step of our algorithm (in the experiments in Section 4) sets the weights of the input ($I$), hidden ($H_{i}$, $11\leq i\leq n-$ 1), and output ($O$) layers of the network as follows:
@@
$H_{i}(x)=\begin{bmatrix}1&\pm[S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})]&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&0\\ 0&S_{i}/a_{i}&-S_{i}/(a_{i}-a_{i}^{2})&-S_{i}a_{i+1}\\ 0&0&0&S_{i}\end{bmatrix}\times ReLU(H_{i-1}(x))$
$O(x)=\begin{bmatrix}1&S_{n}/a_{n}&-S_{n}/(a_{n}-a_{n}^{2})&0\end{bmatrix}\times ReLU(H_{n-1}(x)),$
where $S_{i}$ can either be chosen independently or chosen based on $A$. In the latter case, Equation 3 gives $S_{i}=s_{i}/s_{i-1}=(1-a_{i})a_{i+1}$. For this version, one must decide whether to add or subtract triangle waves from the sum neuron based on whether the target function is convex or concave, respectively. The version of the algorithm used in Section 5 onward does not have this stipulation and always follows Equation 3. Additionally, it expands the input domain to $[-1,1]$ and flips the triangles upside down to compose “v”-shaped functions instead:
$I(x)=\begin{bmatrix}0\\ \frac{-0.5}{a_{0}}\\ \frac{0.5}{1-a_{0}}\\ 0\end{bmatrix}x+\begin{bmatrix}1\\ 1-\frac{0.5}{a_{0}}\\ 1-\frac{0.5}{1-a_{0}}\\ 1\end{bmatrix}$
$H_{i}(x)=\begin{bmatrix}1&1&1&-1\\ 0&a_{i+1}-1&a_{i+1}-1&a_{i}(1-a_{i+1})\\ 0&\frac{a_{i}(1-a_{i+1})}{1-a_{i}}&\frac{a_{i}(1-a_{i+1})}{1-a_{i}}&\frac{-a_{i}^{2}(1-a_{i+1})}{1-a_{i}}\\ 0&0&0&a_{i}(1-a_{i+1})\end{bmatrix}\times ReLU(H_{i-1}(x))$
$O(x)=\begin{bmatrix}1&1&1&-1\end{bmatrix}\times ReLU(H_{n-1}(x)).$
This assignment of $I$, $H$, and $O$ is used in each iteration of the pretraining algorithm to set the weights of each layer before the forward pass. The backward pass can then propagate the gradients back through this weight setting procedure to update the vector $A$ containing the triangle peaks.
After pretraining, the weights can then be initialized once more based on the learned vector $A$, and then updated directly using regular gradient descent in a second round of optimization. Full pseudocode is listed in Algorithm 1.
@@ A.2 Terminology
Before presenting further formal mathematical details of our method, we first briefly review a few pieces of basic ReLU network terminology used in the paper. The reader is referred to Chmielewski-Anders (2020) for an excellent alternative presentation of these terms.
@@ A.3 Necessary Conditions for Differentiability
$F(x)=\sum_{i=0}^{\infty}s_{i}W_{i}(x).$
The goal of this section is to show how to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined and continuous on all of $[0,1]$. We begin by assuming that
$F^{\prime}(x)=\sum_{i=0}^{\infty}s_{i}W^{\prime}_{i}(x).$
We will see that the resulting choice of $s_{i}$ ensures uniform convergence of the derivative terms, so that the derivative of the infinite sum is indeed infinite sum of the derivatives. Fortunately, the left and right derivatives $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ already exist everywhere, since each bend in each $W_{i}$ (where the full derivative of $F$ is undefined) is assigned the slope of the line segment to its left or right, respectively. The $s_{i}$ scaling values will have to be chosen appropriately so that $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ are equal for all bend points.
@@
Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$. We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends. $B_{i}:=P_{i}\cup V_{i}$. $f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$. The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network. This odd split around layer $i$ makes the proofs cleaner.
Figure 10 highlights some important properties about composing triangle functions. Peaks alternate with valleys. Peak locations in one layer become valleys in the next. Valleys in one layer remain valleys in all future layers since 0 is a fixed point of each $T_{i}$. To produce $W_{i}$, each line segment of $W_{i-1}$ becomes a dilated copy of $T_{i}$. Each triangle function has two distinct slopes, $1/a_{i}$ and $-1/(1-a_{i})$, which are dilated by the chain rule during the composition. On negative slopes of $W_{i-1}$, the input to layer $i$ is reversed, so those copies of $T_{i}$ are reflected. Due to the reflection, the slopes of $W_{i}$ on each side of a peak or valley are proportional. Alternatively, one could consider that on each side of a peak in $W_{i-1}$, there is a neighborhood of points that are greater than $a_{i}$, and are composed with the same line segment of $T_{i}$ that has slope $-1/(1-a_{i})$. Either way, it is important to note that the slopes on each side of a bend scale identically during each composition.
@@
Let $x_{k}$ be some point in $P_{i}$, and let $k$ be its index in any list it appears in. To calculate the value of $E_{+}^{\prime}(x_{k})=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x_{k})$, we will have to find the slope of the linear intervals to the immediate right of $x_{k}$ for all $W_{i}$. We will use $R_{x}$ to represent $W_{i+}^{\prime}(x_{k})$. The first term in the sum will be $R_{x}s_{i}$. Since the derivatives of composed functions will multiply from the chain rule, so the value of the next term is $W^{\prime}_{i+1}(x_{k})=T^{\prime}_{i+1}(W_{i}(x_{k}))R_{x}$. $T_{i+1}$ has two linear segments, giving two slope possibilities to multiply by. The correct one to choose is $-1/(1-a_{i+1})$ because it is “active” around $x_{k}$ ($x_{k}$ is a peak of $W_{i}$, so $W_{i}(x_{k})>a_{i+1}$ for $x\in(B_{i+1}[k-1],B_{i+1}[k+1])$). This gives $W^{\prime}_{i+1}(x_{k})=-R_{x}\frac{s_{i+1}}{(1-a_{i+1})}$. Note that the second term has the opposite sign as the first.
For all remaining terms, since $x_{k}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$. For $x\in(B_{j+1}[k-1],B_{j+1}[k+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the slope $1/a_{j+1}$. Since this slope is positive, every remaining term continues to have the opposite sign as the first term. Summing up all the terms with the coefficients $s_{i}$, and factoring out $R_{x}$ will yield the desired formula. Note that this same derivation applies to the left sided derivatives as well because the “active” slopes of $T_{i+1}(W_{i})$ are all the same whether a bend in $W_{i}$ is approached from the left or the right. The initial slope constant $L_{x}$ will just be different. ∎
@@
$F^{\prime}(x)=f^{\prime}_{i}(x)=\sum_{j=0}^{i-1}s_{j}W^{\prime}_{j}(x).$ (5)
@@
We now prove our main theorem, which shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting approximation converges to a continuously differentiable function. The idea of the proof is that much of the formula for $E^{\prime}(x)$ will be shared between two successive generations of peaks. Once they are both valleys, they will be treated the same by the remaining compositions, so the sizes of their remaining discontinuities will need to be proportional. {mytheorem}3.1 $F^{\prime}(x)$ is continuous on $[0,1]$ only if the scaling coefficients are selected based on $a_{i}$ according to:
$s_{i+1}=s_{i}(1-a_{i+1})a_{i+2}.$
@@
Figure 11: The derivatives of the first few stages of approximation. Notice that each time a constant segment “splits,” the two neighboring segments adjacent to the split monotonically converge back to the original value (marked with a colored point corresponding to the step it originates from).